What percent of 45 is 7.2?
Answer by stanbon(75887)

What percent of 45 is 7.2?
---------------------------
45 is the larger amount.
7.2 is the smaller amount.
----
100 is all the percent one can have
x is a part of 100
---------------------------
x/100 = 7.2/45
----
x = 100[7.2/45]
----
x = 720/45 = 16
----
Ans: 16%
============
Checking: 16%*45 = (16/100)*45 = (16*45)/100 = 720/100 = 7.2
================
Cheers,
Stan H.
==============